# Part I: General framework, application to Hartree-Fock and DFT

### v1

@@ Abstract
Computing excited states of many-body quantum Hamiltonians is a fundamental challenge in computational physics and chemistry, with state-of-the-art methods broadly classified into variational (critical point search) and linear response approaches. This article introduces a unified framework based on the Kähler manifold formalism, which naturally accommodates both strategies for a wide range of variational models, including Hartree-Fock, CASSCF, Full CI, and adiabatic TDDFT. In particular, this formalism leads to a systematic and straightforward way to obtain the final equations of linear response theory for nonlinear models, which provides, in the case of mean-field models (Hartree-Fock and DFT), a simple alternative to Casida’s derivation. We detail the mathematical structure of Hamiltonian dynamics on Kähler manifolds, establish connections to standard quantum chemistry equations, and provide theoretical and numerical comparisons of excitation energy computation schemes at the Hartree-Fock level.
@@ 1 Introduction
Computing excited states of many-body quantum Hamiltonians is one of the biggest challenges in computational physics and chemistry 1.
@@
    variational or bivariational methods: computation of critical points of energy functionals on differentiable manifolds or algebraic varieties. This category encompasses state-specific and state-averaged methods 2, 3, 4, 5, 6, 7, 8;
@@
    methods based on linear response theory. Linear-response Hartree-Fock 9, 10, 11, 12, 13, 14, 15(LR-HF), Complete Active Space Self-Consistent Field 16, 17, 18, 19 (LR-CASSCF), time-dependent density functional theory 20, 21, 22, 23, 24, 25 (LR-TDDFT), and equation-of-motion coupled cluster 26, 27, 28, 29, 30, 31, 32, 33, 34 (EOM-CC) fall into this category.
All these approaches coincide for the reference electronic Schrödinger equation set on the continuous position space $\bigwedge^{N}L^{2}({\mathbb{R}}^{3};{\mathbb{C}}^{2})$, as well as for the Full Configuration Interaction (FCI) approximation 35, 36, but differ when further approximations are made. This article aims at providing a unified and compact framework covering critical point search (CP) and linear response theory (LR) from a theoretical viewpoint.
We will use the Kähler manifold formalism allowing one to conveniently formulate CP and LR for any variational model, including any variant of HF (restricted, unrestricted, general, restricted open-shell), CASSCF, and FCI methods, in a unified and compact framework. It also applies to adiabatic TDDFT 37, 38, as this method has an intrinsic Hamiltonian Kähler structure, although it is not a variational approximation of the reference electronic Schödinger equation. We will concentrate on the FCI and HF theories in this study. The cases of more advanced variational approximations such as CASSCF 39 and bivariational approximations such as CC 40, 41, 42 will be dealt with in a future work.
@@
    finding critical points $x_{0},x_{1},x_{2},\ldots$ of $\mathcal{E}$ on $\mathcal{M}$ that can be identified with approximate ground and excited states, labelled in such a way that $\mathcal{E}(x_{0})\leq\mathcal{E}(x_{1})\leq\mathcal{E}(x_{2})\leq\cdots$;
@@
Step 1 is quite involved. Indeed, in contrast with linear models such as FCI, not all critical points of nonlinear models such as HF or CASSCF are suitable approximations of ground or excited states of the reference electronic Schrödinger equation. In general, the energy functional $\mathcal{E}$ can have many spurious critical points on $\mathcal{M}$. Worse still, there can be mismatch between the $k$-th excited states and the critical points with a Morse index $k$, i.e., critical points where the Riemannian Hessian has exactly $k$ negative eigenvalues. This phenomenon is illustrated on a 2D toy model in Fig. 1, as well as on an H4 molecule in Section 5.2. In Ref. 43, Lewin introduced a nonlinear min-max characterization of physically-relevant critical points for CASSCF with no internal orbitals. In Ref. 44, an algorithm for computing CASSCF approximations of first excited states based on this characterization was proposed and tested on an H2 molecule.
Figure 1: Contour of a 2D function $E:(0,1)\times(0,1)\to{\mathbb{R}}$ with one global minimizer ($A$), one local, non-global minimizer ($A^{\prime}$) and one saddle point with Morse index 1 ($B$). The restriction $\mathcal{E}$ of this function to the one-dimensional manifold $\mathcal{M}$ (indicated by the orange curve) has 3 global minimizers ($a_{1}$, $a_{2}$, $a^{\prime}$), 0 local, non-global, minimizers, and 2 critical points ($b$, $b^{\prime}$) with a Morse index 1, which are also local maxima since $\mathcal{M}$ is one-dimensional. The points $a_{1}$ and $a_{2}$ can be considered as reasonable approximations of $A$, $a^{\prime}$ as a reasonable approximation of $A^{\prime}$, $b$ as a reasonable approximation of $B$. The point $b^{\prime}$ is a spurious critical point, which does not approximate well any saddle point of $E$.
@@
This approach provides a systematic and conceptually simple way to derive linear-response approximations to excitation energies for any variational model with Kähler structure.
@@
We also show that, denoting by $E_{0}$ the ground-state energy of $\widehat{H}$, by $\Psi_{\star}$ a normalized ground state, and by $\widehat{\Gamma}_{\star}=\ket{\Psi_{\star}}\bra{\Psi_{\star}}$ |, the symplectic eigenvalues of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(\Gamma_{\star})$ coincide with the eigenvalues of $(\widehat{H}-E_{0})|_{\Psi_{\star}^{\perp}}.$ It follows that, as expected,
@@
where $E_{k}$ is the energy of the $k$-th FCI excited state. In Section 3.4, we consider the case of mean-field models, including HF and adiabatic TDDFT. In this case, $n=N_{\rm b}$, $r=N$, $\gamma\in{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}$ represents the one-body reduced density matrix (1-RDM), and the energy functional ${\mathcal{E}}$ is nonlinear. Eq. (1) is then equivalent to the time-dependent HF (TDHF) or adiabatic TDDFT equations, Eq. (2) to the usual HF or Kohn-Sham equations, and Eq. (5) leads to a straightforward derivation of the final equation providing single excitation energies. Recall that this equation is usually inferred from Casida’s equations of LR-TDHF or LR-TDDFT 23. In Section 4, we make use of the compact formalisms (1)-(5) to provide a theoretical comparison of the electronic excitation energy computation methods based on CP on the one-hand and LR on the other hand, at the UHF level of theory in the weakly interacting regime. More precisely, we consider a family of energy functionals of the form
$\mathcal{E}(\lambda,\gamma):={\rm Tr}(h\gamma)+\frac{\lambda}{2}{\rm Tr}(G(\gamma)\gamma)\quad\text{on}\quad\mathcal{M}={\rm Gr}_{{\mathbb{C}}}(N,N_{\rm b}),$
where $h\in{\mathbb{C}}^{N_{\rm b}\times N_{\rm b}}_{\rm herm}$ and $G:{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}\to{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}$ combines the Coulomb and exchange terms (defined later). We compare the UHF electronic excitation energies $\omega^{\rm HF,CP}_{k}(\lambda)$ and $\omega^{\rm HF,LR}_{k}(\lambda)$ obtained by CP and LR respectively for small values of $\lambda$. In Section 5, we report numerical results supporting our theoretical analysis.
@@ 2 Hamiltonian dynamics on Kähler manifolds
In this section, we intend to provide a brief introduction to Kähler manifolds and associated Hamiltonian dynamics. We refer interested readers to e.g., the monographs  45, 46, 47.
@@
It is easy to check that, for any $u$, $v\in T_{x}\mathcal{M}_{{\mathbb{R}}}$,
@@
$d\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to T^{*}\mathcal{M}_{\mathbb{R}}\qquad\mbox{and}\qquad{\rm grad}_{\mathcal{M}}\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to T\mathcal{M}_{\mathbb{R}}$
@@
The equilibrium points (steady states) satisfy
${{\rm grad}_{\mathcal{M}}\mathcal{E}(x_{\star})=0}$ (9)
and the LR around an equilibrium point $x_{\star}$ consists in studying the linearized dynamics around $x_{\star}$ given by
@@
Eq. (10) is a Hamiltonian dynamics on the symplectic space $(T_{x_{\star}}\mathcal{M}_{\mathbb{R}},J_{x_{\star}})$ with quadratic Hamiltonian
@@
Let $(u_{1},\ldots,u_{n})$ be an orthonormal basis of $T_{x_{\star}}\mathcal{M}$ for $\langle\bullet,\bullet\rangle_{x_{\star}}$, i.e., a family of $n$ vectors of $T_{x_{\star}}\mathcal{M}$ such that
@@
and $\mathcal{X}:={\rm Span}_{\mathbb{R}}(u_{1},\ldots,u_{n})$ the $n$-dimensional real-vector space spanned by the $u_{j}$’s. We have
@@
and we infer from Eq. (7) that in the basis $(u_{1},\ldots,u_{n},{J_{x_{\star}}u_{1},\ldots,J_{x_{\star}}u_{n}})$ of $T_{x_{\star}}\mathcal{M}_{\mathbb{R}}$, $g_{x_{\star}}$ is represented by the identity matrix $I_{2n}$, $J_{x_{\star}}$ by the skew symmetric matrix
$J:=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}.$
@@
$\mathfrak{H}_{\star}:=\left(\begin{array}[]{cc}H^{\rm rr}&H^{\rm ri}\\ H^{\rm ir}&H^{\rm ii}\end{array}\right)\in{\mathbb{R}}^{2n\times 2n}_{\rm sym}$ (11)
@@
$[H^{\rm ri}]_{jj^{\prime}}$ $:=g_{x_{\star}}(u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})(J_{x_{\star}}u_{j^{\prime}})),$ (13)
$[H^{\rm ir}]_{jj^{\prime}}$ $:=g_{x_{\star}}(J_{x_{\star}}u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})u_{j^{\prime}}),$ (14)
$[H^{\rm ii}]_{jj^{\prime}}$ $:=g_{x_{\star}}(J_{x_{\star}}u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})(J_{x_{\star}}u_{j^{\prime}})).$ (15)
If $x_{\star}$ is a nondegenerate local minimizer of $\mathcal{E}$ on $\mathcal{M}_{\mathbb{R}}$, then the matrix $\mathfrak{H}_{\star}$ is real symmetric positive definite (SPD). According to Williamson’s theorem 48, there exists a symplectic matrix $S\in{\mathbb{R}}^{2n\times 2n}$ (i.e., a matrix in ${\mathbb{R}}^{2n\times 2n}$ such that $S^{T}JS=J$) such that
$S^{T}\mathfrak{H}_{\star}S=\left(\begin{array}[]{cc}\Omega&0\\ 0&\Omega\end{array}\right)\quad\mbox{with}\quad\Omega:=\mbox{diag}(\omega_{1}^{\rm LR},\ldots,\omega_{n}^{\rm LR}),\quad 0<\omega_{1}^{\rm LR}\leq\cdots\leq\omega_{n}^{\rm LR}.$ (16)
The $\omega_{j}^{\rm LR}$’s are called the symplectic eigenvalues of $\mathfrak{H}_{\star}$. It is easy to show that these quantities are independent of the chosen orthonormal basis $(u_{1},\ldots,u_{n})$ of $T_{x_{\star}}\mathcal{M}$. They are therefore intrinsic and are called the symplectic eigenvalues of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})$.
Denoting by $Y(t)\in{\mathbb{R}}^{2n}$ the column vector containing the coefficients of $y(t)$ in the basis $(u_{1},\ldots,u_{n},J_{x_{\star}}u_{1},\ldots,J_{x_{\star}}u_{n})$ and by
@@
The angular frequencies $\omega_{1},\ldots,\omega_{n}$ therefore represent the excitation frequencies of the system in the linear response regime.
@@
$\mathfrak{H}_{\star}=\left(\begin{array}[]{cc}H^{\rm rr}&0\\ 0&H^{\rm ii}\end{array}\right)\in{\mathbb{R}}^{2n\times 2n}_{\rm sym}.$ (17)
@@
$S^{T}\mathfrak{H}_{\star}S=\left(\begin{array}[]{cc}\widetilde{\Omega}&0\\ 0&\widetilde{\Omega}\end{array}\right),\quad\mbox{with}\quad\widetilde{\Omega}=\left((H^{\rm rr})^{1/2}H^{\rm ii}(H^{\rm rr})^{1/2}\right)^{1/2}.$ (18)
Diagonalizing $\widetilde{\Omega}$ readily provides the symplectic eigenvalues of $\mathfrak{H}_{\star}$.

### v2

@@ Abstract
Computing excited states of many-body quantum Hamiltonians is a fundamental challenge in computational physics and chemistry, with state-of-the-art methods broadly classified into variational (critical point search) and linear response approaches. The Kähler manifold formalism provides a uniform framework which naturally accommodates both strategies for a wide range of variational models, including Hartree-Fock, CASSCF, Full CI, and adiabatic TDDFT. In particular, this formalism leads to a systematic and straightforward way to obtain the final equations of linear response theory for nonlinear models, which provides, in the case of mean-field models (Hartree-Fock and DFT), a simple alternative to Casida’s derivation. We detail the mathematical structure of Hamiltonian dynamics on Kähler manifolds, establish connections to standard quantum chemistry equations, and provide theoretical and numerical comparisons of excitation energy computation schemes at the Hartree-Fock level.
@@ 1 Introduction
Computing excited states of many-body quantum Hamiltonians is one of the biggest challenges in computational physics and chemistry.[1]
@@
    variational or bivariational methods: computation of critical points of energy functionals on differentiable manifolds or algebraic varieties. This category encompasses state-specific and state-averaged methods;[2, 3, 4, 5, 6, 7, 8]
@@
    methods based on linear response theory. Linear-response Hartree-Fock[9, 10, 11, 12, 13, 14, 15](LR-HF), Complete Active Space Self-Consistent Field[16, 17, 18, 19] (LR-CASSCF), time-dependent density functional theory[20, 21, 22, 23, 24, 25] (LR-TDDFT), and equation-of-motion coupled cluster[26, 27, 28, 29, 30, 31, 32, 33, 34] (EOM-CC) fall into this category.
All these approaches coincide for the reference electronic Schrödinger equation set on the continuous position space $\bigwedge^{N}L^{2}({\mathbb{R}}^{3};{\mathbb{C}}^{2})$, as well as for the Full Configuration Interaction (FCI) approximation,[35, 36] but differ when further approximations are made. This contribution aims at analyzing and comparing these methods from a theoretical viewpoint.
For this purpose, we will use the Kähler manifold formalism, allowing one to conveniently formulate both critical point search (CP) and linear response theory (LR) for any variational model,[37] including any variant of HF [38, 39, 40, 41] (restricted, unrestricted, general, restricted open-shell), CASSCF,[42] and FCI[35, 36] methods, in a unified and compact framework. It also applies to adiabatic TDDFT,[43, 44] as this method has an intrinsic Hamiltonian Kähler structure, although it is not a variational approximation of the reference electronic Schrödinger equation. Although the Kähler manifold formalism was previously used in the physics literature to study infinite-dimensional Hamiltonian systems, including the time-dependent Schrödinger equation,[45] or to compute excitation modes of mean-field models (normal mode analysis),[37] it is very little used in quantum chemistry. The only quantum chemistry articles we are aware of following a similar approach are the works by Arponen[46, 47] on Coupled-Cluster theory, and the recent work by Kvaal and coworkers[48] aiming at better understanding the structure of approximation methods for time-dependent quantum dynamics.
We will concentrate on the FCI and HF theories in Part I of this contribution. The cases of more advanced variational approximations such as CASSCF, DMRG, and bivariational approximations such as CC,[49, 50, 51] will be dealt with in Parts II and III.
@@
    finding critical points $x_{0},x_{1},x_{2},\ldots$ of $\mathcal{E}$ on $\mathcal{M}$ that can be identified as approximate ground and excited states, labelled in such a way that $\mathcal{E}(x_{0})\leq\mathcal{E}(x_{1})\leq\mathcal{E}(x_{2})\leq\cdots$;
@@
Step 1 is quite involved. Indeed, in contrast with linear models such as FCI, not all critical points of nonlinear models such as HF or CASSCF are suitable approximations of ground or excited states of the reference electronic Schrödinger equation. In general, the energy functional $\mathcal{E}$ can have many spurious critical points on $\mathcal{M}$. Worse still, there can be mismatch between the $k$-th excited states and the critical points with a Morse index $k$, i.e., critical points where the Riemannian Hessian has exactly $k$ negative eigenvalues.[52] This phenomenon is illustrated on a 2D toy model in Fig. 1, as well as on an H4 molecule in Section 5.2. In Ref. [53], Lewin introduced a nonlinear min-max characterization of physically-relevant critical points for CASSCF with no internal orbitals. In Ref. [54], an algorithm for computing CASSCF approximations of first excited states based on this characterization was proposed and tested on an H2 molecule. We will elaborate further on this characterization in Part II.
Figure 1: Contour of a 2D function $E:(0,1)\times(0,1)\to{\mathbb{R}}$ with one global minimizer (AA), one local, non-global minimizer ($A^{\prime}$) and one saddle point with Morse index 1 ($B$). The restriction $\mathcal{E}$ of this function to the one-dimensional manifold $\mathcal{M}$ (indicated by the orange curve) has 3 global minimizers ($a_{1}$, $_{2}, aa^{\prime}$), 0 local, non-global, minimizers, and 2 critical points ($b$, $b^{\prime}$) with a Morse index 1, which are also local maxima since $\mathcal{M}$ is one-dimensional. The points $a_{1}$ and $a_{2}$ can be considered as reasonable approximations of $A$, $a^{\prime}$ as a reasonable approximation of $A^{\prime}$, $b$ as a reasonable approximation of $B$. The point $b^{\prime}$ is a spurious critical point, which does not approximate well any saddle point of $E$.
@@
This approach provides a systematic and conceptually simple way to derive linear-response approximations to excitation energies for any variational model with Kähler structure. To our knowledge, this approach was first used by Rowe and coworkers[37] in the context of mean-field models.
@@
We also show that, denoting by $E_{0}$ the ground-state energy of $\widehat{H}$, by $\Psi_{\star}$ a normalized ground state, and by $|\widehat{\Gamma}_{\star}=\ket{\Psi_{\star}}\bra{\Psi_{\star}}$, the symplectic eigenvalues of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(\Gamma_{\star})$ coincide with the eigenvalues of $(\widehat{H}-E_{0})|_{\Psi_{\star}^{\perp}}.$ It follows that, as expected,
@@
where $E_{k}$ is the energy of the $k$-th FCI excited state. In Section 3.4, we consider the case of mean-field models, including HF and adiabatic TDDFT. In this case, $n=N_{\rm b}$ is the size of the basis set, $r=N$, $\gamma\in{\mathbb{C}}_{\rm herm}^{N_{\rm b}\times N_{\rm b}}$ represents the one-body reduced density matrix (1-RDM), and the energy functional $\mathcal{E}$} is nonlinear. Eq. (1) is then equivalent to the time-dependent HF (TDHF) or adiabatic TDDFT equations, Eq. (2) to the usual HF or Kohn-Sham equations, and Eq. (5) leads to a straightforward derivation of the final equation providing single excitation energies. Recall that this equation is usually inferred from Casida’s equations of LR-TDHF or LR-TDDFT.[23] In Section 4, we make use of the compact formalisms (1)-(5) to provide a theoretical comparison of the electronic excitation energy computation methods based on CP on the one-hand and LR on the other hand, at the unrestricted Hartree-Fock (UHF) [55, 56, 57, 58] level of theory in the weakly interacting regime. More precisely, we consider a family of energy functionals of the form
$\mathcal{E}^{\rm UHF}_{\eta}(\gamma_{\alpha},\gamma_{\beta}):={\rm Tr}(h\gamma_{\alpha})+{\rm Tr}(h\gamma_{\beta})+\eta{\mathcal{E}}_{\rm Hx}^{\rm UHF}(\gamma_{\alpha},\gamma_{\beta}),$
on the manifold $\mathcal{M}={\rm Gr}_{\mathbb{C}}(N_{\alpha},N_{\rm b})\times{\rm Gr}_{\mathbb{C}}(N_{\beta},N_{\rm b})$, where $N_{\alpha}$ and $N_{\beta}$ are the numbers of spin-up and spin-down electrons respectively, $h\in{\mathbb{C}}^{N_{\rm b}\times N_{\rm b}}_{\rm herm}$ the one-electron Hamiltonian matrix, ${\mathcal{E}}^{\rm UHF}_{\rm Hx}$ the Hartree plus exchange UHF energy functional and $\eta$ the coupling parameter. We compare the UHF electronic excitation energies $\omega^{\rm UHF,CP}_{k}(\eta)$ and $\omega^{\rm UHF,LR}_{k}(\eta)$ obtained by CP and LR respectively for small values of $\eta$. In Section 5, we report numerical results supporting our theoretical analysis.
@@ 2 Hamiltonian dynamics on Kähler manifolds
In this section, we intend to provide a brief introduction to Kähler manifolds and associated Hamiltonian dynamics. We refer interested readers to, e.g., the monographs [59, 60, 61].
@@
It is easy to check that,
  • the bilinear forms $g_{x}$ and $\omega_{x}$ are respectively real-symmetric and real-antisymmetric on $T_{x}\mathcal{M}_{\mathbb{R}}$;
  • $J_{x}$ is invertible, squares to ${\rm Id}_{T_{x}\mathcal{M}_{{\mathbb{R}}}}$, and transforms any vector $u\in T_{x}\mathcal{M}_{\mathbb{R}}$ into a vector $J_{x}u\in T_{x}\mathcal{M}_{\mathbb{R}}$ orthogonal to $u$:
    $g_{x}(u,J_{x}u)=\omega_{x}(u,J_{x}^{2}u)=\omega_{x}(u,-u)=-\omega_{x}(u,u)=0;$ (7)
  • for any $u$, $v\in T_{x}\mathcal{M}_{{\mathbb{R}}}$,
@@
$d\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to(T\mathcal{M}_{\mathbb{R}})^{*}\qquad\mbox{and}\qquad{\rm grad}_{\mathcal{M}}\mathcal{E}:\mathcal{M}_{\mathbb{R}}\to T\mathcal{M}_{\mathbb{R}}$
@@
A geometric interpretation of the above dynamics is sketched in Fig. 2. It follows from Eq. (7) that at each time $t$, the velocity vector $\frac{dx}{dt}(t)$ is orthogonal to the Riemannian gradient of the energy functional $\mathcal{E}$ at point $x(t)$. As a consequence, the energy $\mathcal{E}$ is a conserved quantity.
Figure 2: Geometric interpretation of a trajectory $t\mapsto x(t)$ of the Hamiltonian dynamics (9) on a Kähler manifold $\mathcal{M}$.
Before proceeding further, let us point out that in the simple case of the flat space $\mathcal{M}={\mathbb{C}}^{n}$ endowed with its natural complex Hermitian form $\langle\bm{z},\bm{z}^{\prime}\rangle=\bm{z}^{*}\bm{z}^{\prime}$, the metric, symplectic form, and symplectic operator are respectively given by
$g_{\bm{z}_{0}}(\bm{z},\bm{z}^{\prime})={\rm Re}(\bm{z}^{*}\bm{z}^{\prime}),\qquad\omega_{\bm{z}_{0}}(\bm{z},\bm{z}^{\prime})={\rm Im}(\bm{z}^{*}\bm{z}^{\prime}),\qquad J_{\bm{z}_{0}}\bm{z}=i\bm{z}.$
Decomposing any vector $\bm{z}\in{\mathbb{C}}^{n}$ into its real and imaginary parts as
$\bm{z}=\bm{q}+i\bm{p}\quad\mbox{with}\quad\bm{q},\bm{p}\in{\mathbb{R}}^{n},$
and introducing the classical Hamiltonian $H_{\rm cl}$ defined on the phase space ${\mathbb{R}}^{n}\times{\mathbb{R}}^{n}$ by
$H_{\rm cl}(\bm{q},\bm{p})=\mathcal{E}(\bm{q}+i\bm{p}),$
the dynamics
$\frac{d\bm{z}}{dt}(t)=J_{\bm{z}(t)}{\rm grad}_{\mathcal{M}}\mathcal{E}(\bm{z}(t))$
is equivalent to the classical Hamiltonian dynamics
$\left\{\begin{array}[]{l}\displaystyle\frac{d\bm{q}}{dt}(t)=\frac{\partial H_{\rm cl}}{\partial\bm{p}}(\bm{q}(t),\bm{p}(t)),\\ \\ \displaystyle\frac{d\bm{p}}{dt}(t)=-\frac{\partial H_{\rm cl}}{\partial\bm{q}}(\bm{q}(t),\bm{p}(t)).\end{array}\right.$. (10)
The dynamics (9) can therefore be interpreted as a generalization of the classical Hamiltonian dynamics (10) on the finite-dimensional phase space ${\mathbb{R}}^{n}\times{\mathbb{R}}^{n}$ to finite or infinite-dimensional Kähler manifolds.
Let us come back to the general case of a generic Kähler manifold. As the operator $J_{x}$ is invertible, the equilibrium points (steady states) of the dynamics (9) are characterized by the equation
${\rm grad}_{\mathcal{M}}\mathcal{E}(x_{\star})=0$}. (11)
The equilibrium points are therefore the critical points of the energy functional $\mathcal{E}$ on the state manifold.
The LR around an equilibrium point $x_{\star}$ consists in studying the linearized dynamics around $x_{\star}$ given by
@@
Eq. (12) is a Hamiltonian dynamics on the flat symplectic space $(T_{x_{\star}}\mathcal{M}_{\mathbb{R}},J_{x_{\star}})$ with a quadratic Hamiltonian
@@
The next step consists in introducing a system of coordinates allowing one to identify $(T_{x_{\star}}\mathcal{M}_{\mathbb{R}},J_{x_{\star}})$ with the usual symplectic space $\big({\mathbb{R}}^{n}\times{\mathbb{R}}^{n},J:=\begin{pmatrix}0&I_{n}\\ -I_{n}&0\end{pmatrix}\big)$. For this purpose, we first consider an orthonormal basis $(u_{1},\ldots,u_{n})$ of $T_{x_{\star}}\mathcal{M}$ for $\langle\bullet,\bullet\rangle_{x_{\star}}$, i.e., a family of $n$ vectors of $T_{x_{\star}}\mathcal{M}$ such that
@@
and $\mathcal{X}:={\rm Span}_{\mathbb{R}}(u_{1},\ldots,u_{n})$ the $n$-dimensional real vector space spanned by the $u_{j}$’s. We have
@@
and we infer from Eq. (8) that in the basis $(u_{1},\ldots,u_{n},{-J_{x_{\star}}u_{1},\ldots,-J_{x_{\star}}u_{n}})$ of $T_{x_{\star}}\mathcal{M}_{\mathbb{R}}$, $g_{x_{\star}}$ is represented by the identity matrix $I_{2n}$, $J_{x_{\star}}$ by the skew symmetric matrix
$J:=\begin{pmatrix}0&I_{n}\\ -I_{n}&0\end{pmatrix}$.
@@
$\mathfrak{H}_{\star}^{\rm LR}:=\left(\begin{array}[]{cc}H^{\rm rr}&H^{\rm ri}\\ H^{\rm ir}&H^{\rm ii}\end{array}\right)\in{\mathbb{R}}^{2n\times 2n}_{\rm sym}$ (13)
@@
$[H^{\rm ri}]_{jj^{\prime}}$ $:=g_{x_{\star}}(u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})(-J_{x_{\star}}u_{j^{\prime}})),$ (15)
$[H^{\rm ir}]_{jj^{\prime}}$ $:=g_{x_{\star}}(-J_{x_{\star}}u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})u_{j^{\prime}}),$ (16)
$[H^{\rm ii}]_{jj^{\prime}}$ $:=g_{x_{\star}}(-J_{x_{\star}}u_{j},{\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})(-J_{x_{\star}}u_{j^{\prime}})).$ (17)
If $x_{\star}$ is a nondegenerate local minimizer of $\mathcal{E}$ on $\mathcal{M}_{\mathbb{R}}$, then the matrix $\mathfrak{H}_{\star}^{\rm LR}$ is real symmetric positive definite (SPD). According to Williamson’s theorem, [62] there exists a symplectic matrix $S\in{\mathbb{R}}^{2n\times 2n}$ (i.e., a matrix in ${\mathbb{R}}^{2n\times 2n}$ such that $S^{T}JS=J$) such that
$S^{T}\mathfrak{H}_{\star}^{\rm LR}S=\left(\begin{array}[]{cc}\Omega&0\\ 0&\Omega\end{array}\right)\quad\mbox{with}\quad\Omega:=\mbox{diag}(\omega_{1}^{\rm LR},\ldots,\omega_{n}^{\rm LR}),\quad 0<\omega_{1}^{\rm LR}\leq\cdots\leq\omega_{n}^{\rm LR}.$ (18)
The $\omega_{j}^{\rm LR}$’s are called the symplectic eigenvalues of $\mathfrak{H}_{\star}^{\rm LR}$. It is easy to show that these quantities are independent of the chosen orthonormal basis $(u_{1},\ldots,u_{n})$ of $T_{x_{\star}}\mathcal{M}$. They are therefore intrinsic and are called the symplectic eigenvalues of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{\star})$.
Denoting by $Y(t)\in{\mathbb{R}}^{2n}$ the column vector containing the coefficients of $y(t)$ in the basis $(u_{1},\ldots,u_{n},-J_{x_{\star}}u_{1},\ldots,-J_{x_{\star}}u_{n})$ and by
@@
The angular frequencies $\omega_{1}^{\rm LR},\ldots,\omega_{n}^{\rm LR}$ therefore represent the excitation frequencies of the system around the local minimizer $x_{\star}$ in the linear response regime.
@@
$\mathfrak{H}_{\star}^{\rm LR}=\left(\begin{array}[]{cc}H^{\rm rr}&0\\ 0&H^{\rm ii}\end{array}\right)\in{\mathbb{R}}^{2n\times 2n}_{\rm sym}.$ (19)
@@
$S^{T}\mathfrak{H}_{\star}^{\rm LR}S=\left(\begin{array}[]{cc}\widetilde{\Omega}&0\\ 0&\widetilde{\Omega}\end{array}\right),\quad\mbox{with}\quad\widetilde{\Omega}=\left((H^{\rm rr})^{1/2}H^{\rm ii}(H^{\rm rr})^{1/2}\right)^{1/2}.$ (20)
Diagonalizing $\widetilde{\Omega}$ readily provides the symplectic eigenvalues of $\mathfrak{H}_{\star}^{\rm LR}$.
The Kähler formalism therefore provides a systematic way to derive linear response excitation energies for any variational model, which can be summarized in the right column of Table 1 below. The key step is the identification of the canonical basis of the tangent space $T_{x_{0}}\mathcal{M}$ at the ground state $x_{0}$. The rest is just straightforward computation of matrix elements using Slater-Condon rules, and application of generic results of linear algebra. The situation is somewhat similar to the computation of truncated CI excited states, where the key step is the choice of the canonical basis of the truncated CI space, composed of the HF ground state $\Phi_{0}$, the single excitations $\Phi_{i}^{a}$, the double excitations $\Phi_{ij}^{ab}$, etc.
Table 1: Steps for computing excitation energies in truncated CI (here CISD as a matter of illustration) and LR for variational models.
CISD LR for variational models
Compute an HF ground state
$\Phi_{0}\in\mathop{\rm argmin}_{\Phi}E^{\rm HF}(\Phi)$
Compute a ground state of the model
$x_{0}\in\mathop{\rm argmin}_{x\in\mathcal{M}}\mathcal{E}(x)$
Identity the CISD subspace canonical basis
$\mathcal{H}^{\rm CISD}=\mbox{Span}(\Phi_{0},\Phi_{i}^{a},\Phi_{ij}^{ab})$
Identify the canonical basis of $T_{x_{0}}\mathcal{M}_{{\mathbb{R}}}$
$T_{x_{0}}\mathcal{M}_{\mathbb{R}}=\mbox{Span}_{\mathbb{R}}(u_{1},\ldots,u_{n},-J_{x_{0}}u_{1},\ldots,-J_{x_{0}}u_{n})$
Build the matrix of $\widehat{H}$ in this basis
$\mathfrak{H}^{\rm CISD}_{0}=\left(\begin{array}[]{ccc}E_{0}^{\rm HF}&\langle\Phi_{0}|\widehat{H}|\Phi_{i}^{a}\rangle&\langle\Phi_{0}|\widehat{H}|\Phi_{ij}^{ab}\rangle\\ &\langle\Phi_{i}^{a}|\widehat{H}|\Phi_{j}^{b}\rangle&\langle\Phi_{i}^{a}|\widehat{H}|\Phi_{jk}^{bc}\rangle\\ &*&\langle\Phi_{ij}^{ab}|\widehat{H}|\Phi_{kl}^{cd}\rangle\end{array}\right)$
Build the matrix of ${\rm Hess}_{\mathcal{M}}\mathcal{E}(x_{0})$ in this basis
$\mathfrak{H}_{0}^{\rm LR}=\left(\begin{array}[]{cc}\mathfrak{h}_{qq}&\mathfrak{h}_{qp}\\ \mathfrak{h}_{qp}^{T}&\mathfrak{h}_{pp}\end{array}\right)$
Compute the eigenvalues of $\mathfrak{H}^{\rm CISD}_{0}$ Compute the symplectic eigenvalues of $\mathfrak{H}^{\rm LR}_{0}$